OpenStudy (anonymous):

The product of two perfect square numbers minus their sum is 4844, find these 2 numbers

OpenStudy (anonymous):

you mean \[a^2 b^2 -(a^2+b^2)=4844\] ?

OpenStudy (anonymous):

use this \[a^2 b^2 -(a^2+b^2)=4844\\a^2 b^2 -a^2-b^2+1=4845\\ a^2(b^2-1)-(b^2-1)=4845\\(a^2-1)(b^2-1)=4845=3*5*17*19\]

OpenStudy (asnaseer):

Neat solution @mukushla

OpenStudy (anonymous):

@asnaseer tnx